jhonyy9 (jhonyy9):

- n,x and y are integers, greater than zero how many solutions (x,y) has the below equation ? x^2 -y^2 = 10^n

jhonyy9 (jhonyy9):

@ganeshie8 what is your opinion about this please ? ty.

jhonyy9 (jhonyy9):

we can rewriting x^2 -y^2 = (x-y)(x+y) and 10^n = 2^n *5^n

OpenStudy (loser66):

I think it is only 1 solution since it is a hyperbola \[\dfrac{x^2}{10^n}-\dfrac{y^2}{10^n}=1\] when y=0, \(x^2=10^n\rightarrow x=\pm 10^{n/2}\) and it is only one solution for the hyperbola.